can someone please help me?
solve each system by the elimination method.
1. 2x-y=2
4x+y=10

Answer by ankor@dixie-net.com(22740)

2x - y = 2
4x + y = 10
----------------adding the equations eliminates y, find x
6x = 12
x = 12/6
x = 2
:
Find y using the 2nd equation, replace x with 2
4(2) + y = 10
8 + y = 10
y = 10 - 8
y = 2
:
Check solutions in the 1st equation
2(2) - 2 = 2
4 - 2 = 2